If you have 2.9 L of gas at standard pressure and temperature, what will be the pressure of the gas be if you decrease the volume to 2.4 L and kept standard pressure?
1
Expert's answer
2020-10-07T07:34:11-0400

As Pressure inversely proportional to Volume.

P1/P2= V2/V1

at STP (pressure = 1 bar, temperature = 273 K)

1/P2=2.4/2.9

P2= 1.208 bar